# Defining subsystems in Hilbert spaces with non-Euclidean metric

### v2

@@ Appendix C Spin tomography
General spin tomography: Given eigenvectors $|m\rangle$ of $\sigma_{z}$ with eigenvalues $m\in\{\dfrac{1}{2},-\dfrac{1}{2}\}$, the probability of obtaining the result $m$ after measuring the spin $\sigma_{z}$ is given by $p(m|\sigma_{z})=\langle m|\rho|m\rangle$. Consider a unitary evolution of this state under $U=(\mathbb{1}-i\sigma_{Y})/\sqrt{2}$, the subsequent measurement of spin in the $\sigma_{3}$ direction, the probability of getting the result $m$, given by $p(m)=\langle m|U^{\dagger}\rho U|m\rangle$ is the probability of getting the result $m$ after a $\sigma_{1}$ measurement on $\rho$. This is because $U|m\rangle$ is the eigenvector of $\sigma_{x}$ with eigenvalue $m$. In general, if one wants to measure the observable $\vec{\sigma}.\vec{n}$ where $\vec{n}=(\sin\theta\cos\phi,\sin\theta\sin\phi,\cos\theta)$, it is equivalent to a $\sigma_{z}$ measurement of the state $U\rho U^{\dagger}$ where $U=e^{-i(\theta/2)\sigma.\hat{n}_{\perp}}$ is unitary with $\hat{n}_{\perp}=(-\sin\phi,\cos\phi,0)$. This is possible due to the following identity: $e^{-i(\theta/2)\sigma.\hat{n}_{\perp}}\sigma_{z}e^{i(\theta/2)\sigma.\hat{n}_{\perp}}=\vec{\sigma}.\vec{n}$  . In the Stern-Gerlach setting, the interaction Hamiltonian is given by $H=-\vec{\mu}.\vec{B}$, where $\vec{\mu}=\gamma\vec{\sigma}$ is the magnetic moment, $\gamma$ is the gyromagnetic factor, and $\vec{B}$ is the magnetic field. The evolution $U$ is carried out by fixing the this evolution is carried out by the magnetic field $\vec{B}=\theta\hat{n}_{\perp}/(\gamma t)$.
Given the data $\{m,p(m|\hat{n}_{i})\}$, one can construct the density matrix using the Born rule. The probabilities are related to the state in the following way: $p(m|n_{3})=\langle m|\rho|m\rangle$, such that $\sum mp(m|n_{3})=tr(\rho\sigma_{3})$ is the expectation value of the $\sigma_{3}$ measurement. Fixing the magnetic field such that $\gamma Bt\vec{\sigma}.\hat{n}=\pi/4\sigma_{1}$  and then passing the particle through a secondary magnetic field in $z-$direction, one measures $\sigma_{2}$. The probabilities are given by $p(m|n_{1})=\langle m|U_{1}\rho U_{1}^{\dagger}|m\rangle$ where $U_{1}=e^{-i\pi/4\sigma_{2}}$. Similarly, for $U_{2}=e^{i\pi/4\sigma_{1}}$, we have $p(m|n_{2})=\langle m|U_{2}\rho U_{2}^{\dagger}|m\rangle$.

### v3

@@ Appendix C Spin tomography
We will outline the method to tomographically construct a single particle spin 1/2 system. Given eigenvectors $|m\rangle$ of $\sigma_{z}$ with eigenvalues $m\in\{\dfrac{1}{2},-\dfrac{1}{2}\}$, the probability of obtaining the result $m$ after measuring the spin $\sigma_{z}$ is given by $p(m|\sigma_{z})=\langle m|\rho|m\rangle$. Consider a unitary evolution of this state under $U=(\mathbb{1}-i\sigma_{Y})/\sqrt{2}$, the subsequent measurement of spin in the $\sigma_{3}$ direction, the probability of getting the result $m$, given by $p(m|\hat{n}_{1})=\langle m|U^{\dagger}\rho U|m\rangle$ is the probability of getting the result $m$ after a $\sigma_{1}$ measurement on $\rho$. This is because $U|m\rangle$ is the eigenvector of $\sigma_{x}$ with eigenvalue $m$. In general, if one wants to measure the observable $\vec{\sigma}.\vec{n}$ where $\vec{n}=(\sin\theta\cos\phi,\sin\theta\sin\phi,\cos\theta)$, it is equivalent to a $\sigma_{z}$ measurement of the state $U\rho U^{\dagger}$ where $U=e^{-i(\theta/2)\sigma.\hat{n}_{\perp}}$ is unitary with $\hat{n}_{\perp}=(-\sin\phi,\cos\phi,0)$. This is possible due to the following identity: $e^{-i(\theta/2)\sigma.\hat{n}_{\perp}}\sigma_{z}e^{i(\theta/2)\sigma.\hat{n}_{\perp}}=\vec{\sigma}.\vec{n}$  .
Figure 2: Spin tomography of a spin 1/2 particle. The filter $F$ lets particles of a certain velocity pass, which are then unitarily rotated by the magnetic field in the x-y plane. As they pass through the second magnetic field in z-direction, the two arms are detected, corresponding to $m=\pm 1/2$ values.
In the stern Gerlach setting, as shown in figure 2, a beam of electrons passes through a Fizeau filter, which allows particles with a specific velocity to pass through. This is required because the unitary rotation $U$ depends on the time for which the particle interacts with the magnetic field $\vec{B}$. The interaction Hamiltonian is given by $H=-\vec{\mu}.\vec{B}$, where $\vec{\mu}=\gamma\vec{\sigma}$ is the magnetic moment, $\gamma$ is the gyromagnetic factor and $\vec{B}$ is the magnetic field. The evolution $U$ is carried out by fixing this evolution is carried out by the magnetic field $\vec{B}=\theta\hat{n}_{\perp}/(\gamma t)$. After this, the particle passes through a magnetic field in $z$ direction which splits the beam into two, interpreted as ±$\pm 1/2$ detection.
Given the data $\{m,p(m|\hat{n}_{i})\}$, one can construct the density matrix using the Born rule. The probabilities are related to the state in the following way: $p(m|n_{3})=\langle m|\rho|m\rangle$, such that $\sum mp(m|n_{3})=\dfrac{1}{2}\operatorname{tr}(\rho\sigma_{3})$ is the expectation value of the $\sigma_{3}$ measurement. Fixing the magnetic field such that $\gamma Bt\vec{\sigma}.\hat{n}=\pi/4\sigma_{1}$  and then passing the particle through a secondary magnetic field in $z-$direction, one measures $\sigma_{2}$. The probabilities are given by $p(m|\hat{n}_{1})=\langle m|U_{1}\rho U_{1}^{\dagger}|m\rangle$ where $U_{1}=e^{-i\pi/4\sigma_{2}}$. Similarly, for $U_{2}=e^{i\pi/4\sigma_{1}}$, we have $p(m|\hat{n}_{2})=\langle m|U_{2}\rho U_{2}^{\dagger}|m\rangle$.